# NASP: Network Slice as a Service Platform for 5G Networks

### v2

@@ 3 Related Work
Cloud-native platforms are likewise paving the way for zero-touch cores. Scotece2023-5gkube presented 5G-Kube, a Kubernetes-native CNF stack that cuts slice instantiation times below 200 ms. Larrea2023 developed CoreKube, a cloud-native mobile core whose stateless, message-focused workers are containerized and orchestrated by Kubernetes. Additionally, zhao2025 combined adaptive RAN slicing with hierarchical DRL, achieving E2E KPIs across RAN, TN, and CN. Moreover, Wyszkowski2024 present a detailed tutorial that addresses a blind spot in slicing research and standardization: the design-time organization of slice and subnet instances. They systematize a taxonomy of slicing entities and formalize the slice/subnet design process, clarifying concepts not uniformly covered in 3GPP or ETSI. Building on this, the authors enumerate the essential information items a complete design must capture and propose a standards-aligned structure to record them. The tutorial decomposes design into fine-grained, ordered activities and offers automation guidelines that evolve from assisted handling to fully unattended, declarative workflows capable of dynamic slicing.
@@
Despite these state of the art advances, three challenges remain: (i) E2E designs that span RAN, TN, and CN within a SBA; (ii) fully closed-loop slice lifecycle management aligned with SLA objectives; and (iii) an operational NSaaS framework that unifies these capabilities. The NASP framework addresses these gaps by (i) translating GSMA business templates into domain-specific descriptors, (ii) orchestrating hierarchical slice instantiation across RAN, TN, and CN via open interfaces, such as Radio Access Network Interface (RN1), Transport Network Interface (TN1), and Core Network Interface (CN1); and (iii) embedding AI-driven closed-loop assurance. Compared with other solutions, NASP integrates native B5G support, run-time optimization, and a demonstrable NSaaS prototype, illustrating a viable path toward zero-touch, ZSM-compliant B5G deployments.
@@ 4 NSaaS Platform Overview
This section introduces NASP, which aligns with concepts from 3GPP, ETSI, and GSMAframeworks to enable E2E slicing. The platform mediates between business intent and network implementation by automatically translating high‑level business templates into NS definitions, enabling multi‑domain resource orchestration and monitoring. In this context, NASP translates GSMA business templates into versioned Network Slice Templatess through a hierarchical deployment model. Business requests collected by the CSMF are transformed into service requirements and forwarded to the NSMF, which assigns slice identifiers (IDs) and delegates domain‑specific tasks. Standardized interfaces (RN1, TN1, and CN1) ensure seamless integration across the RAN, TN, and CN. Automation is achieved using CI/CD and IaC practices, while a closed‑loop monitoring system enforces continuous telemetry‑based SLA compliance.
@@
This NASP module forms the northbound interface through which tenants submit slice requests encoded using the GSMA GST. Incoming service intents first pass through the Requirements micro‑service, where SLA parameters such as latency, throughput, and availability are normalized. The data stream is forwarded to Template Matching, whose rule‑based engine correlates the normalized attributes with a 3GPP‑compliant library of versioned NSSTs, selecting the candidate that minimizes resource footprint while maximizing reuse.
Once a suitable template is chosen, Resource Translations converts logical requirements into concrete descriptors for virtual compute, storage, radio, and transport resources, using mapping rules derived from TMF638; TMF639 and 3gpp.28.541. The translated artifact is forwarded to the Descriptions service, which renders a standards‑compliant Network Slice Descriptor (NSD) in YAML Ain’t Markup Language (YAML) format, directly consumable by the orchestration tier. A GitOps‑driven CI/CD pipeline performs syntax validation, policy checks, and version tagging before committing the artifact to an immutable registry. During onboarding, each GST attribute, such as e2eLatency, serviceAvailability, is persisted as a ServiceSpecification/SlaTarget object via TeleManagement Forum (TMF) 638 (v5.0.0). These targets are version‑controlled alongside the NS descriptors and immediately promoted to TMF 640 (v5.0.0) Service Level Objective (SLO) records so that downstream orchestration and assurance functions can reference a single, immutable source of truth.
@@
The end-to-end workflow, illustrated in Figure 5, begins when the tenant submits a slice request to the CSMF (01); the request is forwarded to the NSMF, which reserves an Single Network Slice Selection Assistance Information (S-NSSAI), retrieves the corresponding slice templates and engages the CN NSSMF to deploy the required CN VNFs under the supervision of the Network Function Virtualization Orchestrator (NFVO) Controller (02–04). Once an AMF endpoint is available, the RAN NSSMF instantiates the next-generation Node B (gNB) components through the RAN Controller so that they attach to the correct AMF (05–07). After both RAN and CN domains are operational, the TN NSSMF instructs the Wide Area Network (WAN) SDN Controller to establish transport connectivity, reports completion to the NSMF and triggers the final activation confirmation that is relayed to the tenant via the CSMF (08–10). All configuration artifacts are stored in the Inventory through idempotent transactions, guaranteeing that repeated executions converge to a consistent slice state, and the final slice configuration is preserved for subsequent monitoring and lifecycle management.
The transformation from high‑level business commitments into enforceable artifacts occurs in four logical phases. (i) Production: GST attributes are normalized and stored as TMF 638 (v5.0.0) SlaTarget objects, each of which is linked, via a mapping table inside the Resource Translations micro‑service, to the corresponding TMF 639 (v5.0.0) resource identifiers, e.g., transport bandwidth pools or radio‑spectrum slices. (ii) Decomposition: the CSMF converts these targets into TMF 640 (v5.0.0) SLO records and distributes them to the NSMF and the domain‑specific NSSMFs, which embed the derived KPI thresholds in their local controllers. (iii) Enforcement: during operation, the Quality‑Assurance loop continuously compares real‑time telemetry with the injected thresholds; any breach raises a TMF 642 (v5.0.0) notification that automatically triggers scaling, relocation, or healing workflows through TMF 645 (v5.0.0) Assurance APIs. (iv) Tenant visibility: current SLA status is exposed through a TMF 645 (v5.0.0) /assurance/v1/monitorings endpoint, and tenants may subscribe to NEF webhooks for proactive breach alerts.
@@
This module supervises every active NS through a non‑real‑time control loop that begins with Telemetry Discovery, which retrieves multilayer observability data (counters, alarms, traces, and logs) via the NEF, RAN monitoring interfaces, and SDN probes. Although transport-domain SDN telemetry (OpenFlow and sFlow Protocol (sFlow)/ IP Flow Information Export (IPFIX)) is not natively covered by 3GPP TS 28 5xx, a lightweight adaptor inside the Data Translation micro-service enriches every record with the S-NSSAI and converts raw counters, such as port-tx_bytes, queue delay, packet-drop events, into their 3GPP-compliant counterparts, defined in 3gpp.28.552. The normalized KPIs are pushed (i) towards the NWDAF for analytics or (ii) as TMF 642 events for fault management. This translation keeps the entire assurance loop 3GPP-compliant while still exploiting the high-resolution visibility provided by SDN probes. The records are normalized and time‑stamped before publication on a common bus. Data Aggregation aligns the heterogeneous streams on the slice ID, such as S‑NSSAI, and enriches them with business KPIs from the catalog. Finally, Data Translation harmonizes vendor‑specific semantics, converting resource counters into utilization metrics, mobility events into experience indicators, and sharing information into isolation scores, while mapping slice intents onto the configuration vocabulary of the Service Management and Orchestration (SMO), NSMF, and domain‑specific NSSMFs.
@@
The NASP control plane exposes three domain‑agnostic southbound interfaces that enable uniform NS lifecycle automation across the RAN, TN, and CN domains. Each interface conveys an intent encoded as a YAML document that augments the 3GPP network‑resource model in TS 28.541 (v19.4.0, Rel‑19) with GST attributes defined by GSMA NG.116 (v10.0). RN1 targets the radio‑access network, TN1 the transport network, and CN1 the mobile core. For example, RN1 submits a RAN Slice Set to the Kubernetes (K8s) API server managing the cloud‑native RAN. The manifest is deployed through Helm releases, Network Attachment Definitions, and accompanying custom resources configuring the required gNB functions. Moreover, TN1 delivers a Path Intent over a RESTful Hypertext Transfer Protocol (HTTP)/2 channel to the SDN controller, translating the request into OpenFlow 1.5 flow‑mod messages. Finally, CN1 forwards a CN Slice Set to the K8s API server that orchestrates the CN. Therefore, the manifest is applied as Helm releases and ancillary resources, including a Data Plane Policy object that configures the user‑plane function datapath.
@@ 5 Prototype Implementation
NASP manipulates IP tables inside Kubernetes pods for TN configuration. Each pod’s network interfaces are dynamically reconfigured to route traffic using VLAN-based segmentation, providing isolation and meeting performance requirements (Figure 7). Distinct VLAN IDs enforce specific QoS policies: a short path prioritizes low latency, while a long path provides redundancy. Although the TN is programmed via ONOS/OpenFlow, we keep a lightweight 802.1Q tag on every packet. Distinct VLAN IDs encode slice-specific QoS policies: a short path prioritizes low latency for URLLC, whereas a longer path provides redundancy for mMTC and Shared slices. Although the TN is programmed via ONOS/OpenFlow, packets retain lightweight 802.1Q tags. These tags act as in-band slice and QoS markers, allowing switches to differentiate flows locally, reduce flow-table pressure, and perform deterministic path selection before any controller interaction. This mechanism preserves layer-2 isolation and remains compatible with legacy Ethernet segments (IEEE8021Q-2018).
@@
Quality Assurance Closed Loop module: Implements telemetry collection (Prometheus exporters), KPI extraction (deployment time, attach latency, availability, resource consumption), and threshold- or rule-based corrective actions such as AMF reconfiguration. Advanced AI-driven analytics (e.g., predictive scaling, anomaly classification) are currently placeholders for future work.
@@
We instantiated four slice scenarios to produce the results discussed later: (i) URLLC, (ii) mMTC (a.k.a., mIoT in later figures), (iii) Shared (an eMBB-type slice that reuses existing control‑plane Network Functions to measure best‑case provisioning), and (iv) non‑3GPP access (offload via N3IWF). Table 2 summarizes the architectural deltas that drive the different deployment and performance outcomes.
@@ 6 Evaluation Methodology
Experiments were executed on Virtual Machines provided by DigitalOcean. The testbed comprises heterogeneous VMs running Ubuntu 22.04 LTS with Linux kernel 5.15.0-88-generic. Control-plane components were deployed on s-4vcpu-8gb instances (4 vCPU, 8 GB RAM), whereas data-plane components used s-8vcpu-16gb instances (8 vCPU, 16 GB RAM). Dedicated VMs were allocated to Open Network Automation Platform (ONAP), Mininet with ONOS, and the Kubernetes cluster. The Kubernetes deployment (v1.28.3) used Calico (Calico) and Multus (Multus) to support multi-tenancy, with autoscaling thresholds configured at 80% for scale-up and 20% for scale-down. Persistent monitoring data were exported to a Prometheus–Grafana stack.
Experiments were conducted on VMs provided by DigitalOcean. The cloud testbed consists of heterogeneous VMs running Ubuntu 22.04 LTS with Linux kernel 5.15.0‑88‑generic. Control‑plane components ran on s‑4vcpu‑8gb instances (4 Virtual CPU (vCPU), 8 GB RAM), whereas data‑plane elements used s‑8vcpu‑16gb instances (8 vCPU, 16 GB RAM). We dedicated a separate VMs to ONAP, Mininet with ONOS, and a Kubernetes cluster. We used Kubernetes v1.28.3 with Calico and Multus to support multi‑tenancy, with autoscaling thresholds of 80% (scale‑up) and 20% (scale‑down). Persistent monitoring data were exported to a Prometheus–Grafana stack, and all timestamps were converted to Coordinated Universal Time (UTC) to avoid clock‑skew artifacts.
In this work, we propose three types of cloud sites for the network performance evaluation: Regional (Central), Metropolitan (Edge), and Internal (Extreme‑Low‑Latency Edge), as defined by gustavo2024evaluating. The Internal site offered limited instance types, so we normalized cost comparisons using pricing from all sites. We emulated the network topology using Mininet and the ONOS controller. Figure 8 shows the long‑range cloud‑network topology, where latency samples were collected from geographically separated regions. Metrics were sampled at 1 Hz over a 72-hour period. A cubic non-linear regression (adjusted $R^{2}_{\text{adj}}=0.93$) was used to summarize average-latency trends over a representative 24-hour cycle.

### v3

@@ 3 Related Work
Cloud-native platforms are likewise paving the way for zero-touch cores. Scotece and others (2023) presented 5G-Kube, a Kubernetes-native Cloud-native Network Function (CNF) stack that cuts slice instantiation times below 200 ms. Larrea et al. (2023) developed CoreKube, a cloud-native mobile core whose stateless, message-focused workers are containerized and orchestrated by Kubernetes. Additionally, Zhao et al. (2025) combined adaptive RAN slicing with hierarchical DRL, achieving E2E KPIs across RAN, TN, and CN. Moreover, Wyszkowski and others (2024) present a detailed tutorial that addresses a blind spot in slicing research and standardization: the design-time organization of slice and subnet instances. They systematize a taxonomy of slicing entities and formalize the slice/subnet design process, clarifying concepts not uniformly covered in 3GPP or ETSI. Building on this, the authors enumerate the essential information items a complete design must capture and propose a standards-aligned structure to record them. The tutorial decomposes design into fine-grained, ordered activities and offers automation guidelines that evolve from assisted handling to fully unattended, declarative workflows capable of dynamic slicing.
@@
Despite these state of the art advances, three challenges remain: (i) E2E designs that span RAN, TN, and CN within a SBA; (ii) fully closed-loop slice lifecycle management aligned with SLA objectives; and (iii) an operational NSaaS framework that unifies these capabilities. The NASP framework addresses these gaps by (i) translating GSMA business templates into domain-specific descriptors, (ii) orchestrating hierarchical slice instantiation across RAN, TN, and CN via open interfaces, such as Radio Access Network Interface (RN1), Transport Network Interface (TN1), and Core Network Interface (CN1); and (iii) embedding AI-driven closed-loop assurance. Compared with other solutions, NASP integrates native B5G support, run-time optimization, and a demonstrable NSaaS prototype, illustrating a viable path toward zero-touch, ZSM-compliant B5G deployments. It is worth noting that while the NASP framework architecture is designed to integrate advanced AI-driven closed-loop assurance, we rate its current B5G/AI Ready status as Medium (M) (see Table 1) to reflect that the current prototype relies mainly on rule-based policies, with full AI integration planned for future work.
@@ 4 NSaaS Platform Overview
This section introduces NASP, which aligns with concepts from 3GPP, ETSI, and GSMAframeworks to enable E2E slicing. The platform mediates between business intent and network implementation by automatically translating high‑level business templates into NS definitions, enabling multi‑domain resource orchestration and monitoring. In this context, NASP translates GSMA business templates into versioned Network Slice Subnet Templates through a hierarchical deployment model. Once initially processed by the Onboard module, business requests are handled by the CSMF, where they are transformed into service requirements and forwarded to the NSMF, which assigns slice identifiers (IDs) and delegates domain‑specific tasks. Standardized interfaces (RN1, TN1, and CN1) ensure seamless integration across the RAN, TN, and CN. Automation is achieved using CI/CD and IaC practices, while a closed‑loop monitoring system enforces continuous telemetry‑based SLA compliance.
@@
This NASP module acts as the single northbound entry point through which tenants submit their slice requests encoded using the GSMA GST, thereby managing the design-time (or preparation) phase. Listing 1 presents a minimal excerpt of a customized eMBB slice request, demonstrating how tenant requirements, such as SLA constraints and domain-specific topological limits, are encoded and submitted.
Listing 1: Example of a custom tenant slice request encoded as a JSON object.
1{ "name": "Custom 5G Network Slice",
2 "NST": {
3 "type": "custom",
4 "Slice Attributes": {
5 "availability": 1,
6 "Supported Data Network": "internet",
7 "SSQ": {
8 "Packet Delay Budget": 0.00012,
9 "Packet Error Rate": 0.0000001,
10 "Maximum Data Burts Volume": 0.001},
11 "UE density": 10000},
12 "resource_description": {
13 "core": { "nfs": [{"name": "amf"}, {"name": "smf"}, {"name": "upf"}] },
14 "ran": { "nfs": [{"name": "ueransim", "type": "gnb", "replicas": 2}] },
15 "tn": { "routes": [{"name": "backhaul"}] }
16 }}}
Incoming service intents first pass through the Requirements micro‑service, where SLA parameters such as latency, throughput, and availability are normalized. The data stream is forwarded to Template Matching, whose rule‑based engine correlates the normalized attributes with a 3GPP‑compliant library of versioned NSSTs, selecting the candidate that minimizes resource footprint while maximizing reuse. If no existing template mapping can satisfy the normalized attributes, the Onboard module rejects the request and returns an error to the tenant. Once the descriptors are prepared and validated, the Onboard module triggers the run-time instantiation phase by forwarding the processed request to the Orchestration module (specifically linking to the CSMF).
Once a suitable template is chosen, Resource Translations converts logical requirements into concrete descriptors for virtual compute, storage, radio, and transport resources, using mapping rules derived from TM Forum (2024a, b) and 3GPP (2025a). The translated artifact is forwarded to the Descriptions service, which renders a standards‑compliant Network Slice Descriptor (NSD) in YAML Ain’t Markup Language (YAML) format, directly consumable by the orchestration tier. A GitOps‑driven CI/CD pipeline performs syntax validation, policy checks, and version tagging before committing the artifact to an immutable registry. During onboarding, each GST attribute, such as e2eLatency, serviceAvailability, is persisted as a ServiceSpecification/SlaTarget object via TeleManagement Forum (TMF) 638 (v5.0.0). These targets are version‑controlled alongside the NS descriptors and immediately promoted to TMF 640 Service Level Objective (SLO) records so that downstream orchestration and assurance functions can reference a single, immutable source of truth.
@@
The end-to-end run-time workflow, illustrated in Figure 5, begins when the tenant’s slice request is forwarded by the Onboard module to the CSMF (01); the request is then forwarded to the NSMF, which reserves an Single Network Slice Selection Assistance Information (S-NSSAI), retrieves the corresponding slice templates and engages the CN NSSMF to deploy the required CN VNFs under the supervision of the Network Function Virtualization Orchestrator (NFVO) Controller (02–04). Once an AMF endpoint is available, the RAN NSSMF instantiates the next-generation Node B (gNB) components through the RAN Controller so that they attach to the correct AMF (05–07). After both RAN and CN domains are operational, the TN NSSMF instructs the Wide Area Network (WAN) SDN Controller to establish transport connectivity, reports completion to the NSMF and triggers the final activation confirmation that is relayed to the tenant via the CSMF (08–10). All configuration artifacts are stored in the Inventory through idempotent transactions, guaranteeing that repeated executions converge to a consistent slice state, and the final slice configuration is preserved for subsequent monitoring and lifecycle management.
The transformation from high‑level business commitments into enforceable artifacts occurs in four logical phases leveraging TMF Open APIs (v5.0.0). (i) Production: GST attributes are normalized and stored as TMF 638 SlaTarget objects, each of which is linked, via a mapping table inside the Resource Translations micro‑service, to the corresponding TMF 639 resource identifiers, e.g., transport bandwidth pools or radio‑spectrum slices. (ii) Decomposition: the CSMF converts these targets into TMF 640 SLO records and distributes them to the NSMF and the domain‑specific NSSMFs, which embed the derived KPI thresholds in their local controllers. (iii) Enforcement: during operation, the Quality‑Assurance loop continuously compares real‑time telemetry with the injected thresholds; any breach raises a TMF 642 notification that automatically triggers scaling, relocation, or healing workflows through TMF 645 Assurance APIs. (iv) Tenant visibility: current SLA status is exposed through a TMF 645 /assurance/v1/monitorings endpoint, and tenants may subscribe to NEF webhooks for proactive breach alerts.
@@
This module supervises every active NS through a non‑real‑time control loop that begins with Telemetry Discovery, which retrieves multilayer observability data (counters, alarms, traces, and logs) via the NEF, RAN monitoring interfaces, and SDN probes. Although transport-domain SDN telemetry (OpenFlow and sFlow Protocol (sFlow)/ IP Flow Information Export (IPFIX)) is not natively covered by 3GPP TS 28 5xx, a lightweight adaptor inside the Data Translation micro-service enriches every record with the S-NSSAI and converts raw counters, such as port-tx_bytes, queue delay, packet-drop events, into their 3GPP-compliant counterparts, defined in 3GPP (2025d). The normalized KPIs are pushed (i) towards the NWDAF for analytics or (ii) as TMF 642 events for fault management. This translation keeps the entire assurance loop 3GPP-compliant while still exploiting the high-resolution visibility provided by SDN probes. The records are normalized and time‑stamped before publication on a common bus. Data Aggregation aligns the heterogeneous streams on the slice ID, such as S‑NSSAI, and enriches them with business KPIs from the catalog. Finally, Data Translation harmonizes vendor‑specific semantics, converting resource counters into utilization metrics, mobility events into experience indicators, and sharing information into isolation scores, while mapping slice intents onto the configuration vocabulary of the Service Management and Orchestration (SMO), NSMF, and domain‑specific NSSMFs.
@@
The NASP control plane exposes three domain‑agnostic southbound interfaces that enable uniform NS lifecycle automation across the RAN, TN, and CN domains. Each interface conveys an intent encoded as a YAML document that augments the 3GPP network‑resource model in TS 28.541 (v19.4.0, Rel‑19) with GST attributes defined by GSMA NG.116 (v10.0). RN1 targets the radio‑access network, TN1 the transport network, and CN1 the mobile core. For example, RN1 submits a RAN Slice Set to the Kubernetes (K8s) API server managing the cloud‑native RAN. The manifest is deployed via Helm releases, Network Attachment Definitions, and accompanying custom resources that configure the required gNB functions. Moreover, TN1 delivers a Path Intent over a RESTful Hypertext Transfer Protocol (HTTP)/2 channel to the SDN controller. Because this intent is technology-agnostic, the controller can translate it into the appropriate southbound protocols based on the underlying network equipment. For instance, it may employ OpenFlow 1.5 flow‑mod messages for fine-grained traffic engineering on virtual switches, or push YANG-modeled configurations via Netconf to manage physical transport routers. Finally, CN1 forwards a CN Slice Set to the K8s API server that orchestrates the CN. Therefore, the manifest is applied as Helm releases and ancillary resources, including a Data Plane Policy object that configures the user‑plane function datapath.
@@ 5 Prototype Implementation
NASP manipulates IP tables inside Kubernetes pods for TN configuration. Each pod’s network interfaces are dynamically reconfigured to route traffic using VLAN-based segmentation, providing isolation and meeting performance requirements (Figure 7). Although the TN is programmed via ONOS/OpenFlow, every packet retains a lightweight 802.1Q tag. Distinct VLAN IDs encode slice-specific QoS policies: a short path prioritizes low latency for URLLC, whereas a longer path provides redundancy for mMTC and Shared slices. These tags act as in-band slice and QoS markers, allowing switches to differentiate flows locally, reduce flow-table pressure, and perform deterministic path selection before any controller interaction. This mechanism preserves layer-2 isolation and remains compatible with legacy Ethernet segments (IEEE, 2018).
@@
Quality Assurance Closed Loop module: Implements telemetry collection (Prometheus exporters), KPI extraction (deployment time, attach latency, availability, resource consumption), and threshold- or rule-based corrective actions such as AMF reconfiguration. Although the architecture includes components for AI-driven analytics, the current implementation relies primarily on rule-based mechanisms, leaving advanced AI capabilities (e.g., predictive scaling, anomaly classification) for future work.
@@
We instantiated four slice scenarios to produce the results discussed later: (i) URLLC, (ii) mMTC (a.k.a., mIoT in later figures), (iii) Shared (an eMBB-type slice that reuses existing control‑plane NFs to measure best‑case provisioning), and (iv) non‑3GPP access (offload via N3IWF). Table 2 summarizes the architectural deltas that drive the different deployment and performance outcomes.
@@ 6 Evaluation Methodology
Experiments were conducted on Virtual Machines provided by DigitalOcean. The cloud testbed comprises heterogeneous VMs running Ubuntu 22.04 LTS with the Linux kernel 5.15.0‑88‑generic. Control‑plane components ran on s‑4vcpu‑8gb instances (4 Virtual CPU (vCPU), 8 GB RAM), whereas data‑plane elements used s‑8vcpu‑16gb instances (8 vCPU, 16 GB RAM). We dedicated a separate VMs to Open Network Automation Platform (ONAP), Mininet with ONOS, and a Kubernetes cluster. We used Kubernetes v1.28.3 with Calico (Calico) and Multus (Multus) to support multi‑tenancy, with autoscaling thresholds of 80% (scale‑up) and 20% (scale‑down). Persistent monitoring data were exported to a Prometheus–Grafana stack, and all timestamps were converted to Coordinated Universal Time (UTC) to avoid clock‑skew artifacts.
In this work, we propose three types of cloud sites for the network performance evaluation: Regional (Central), Metropolitan (Edge), and Internal (Extreme‑Low‑Latency Edge), as defined by Bruno and others (2024). The Internal site offered limited instance types, so we normalized cost comparisons using pricing from all sites. We emulated the network topology using Mininet and the ONOS controller. Figure 8 shows the long‑range cloud‑network topology, where latency samples were collected from geographically separated regions. Metrics were sampled at 1 Hz over a 72-hour period. A cubic non-linear regression (adjusted $R^{2}_{\text{adj}}=0.93$) was used to summarize average-latency trends over a representative 24-hour cycle.